# Addressing the gravitational collapse of a massless scalar field with Physics-Informed Neural Networks

### v1

@@ 7. Define uniform target masses: For each $i$,
In our methodology, a two-dimensional spacetime domain is numerically discretized as $(t,r)\in[0,T]\times[0,R]$, where the parameters $T$ and $R$ denote the maximum values reached in the numerical simulation and must be chosen sufficiently large to encompass all relevant physical information. The resulting grid thus comprises $N_{t}$ points in the temporal dimension and $N_{r}$ in the radial direction. Therefore, the original loss function depends on both time and space, since a loss value is computed at each collocation point. In that sense, the first three terms in the loss equation
@@
The enforcement of restrictions $\left|\mathcal{R}_{\alpha}\right|^{2},\left|\mathcal{R}_{C}\right|^{2}\to 0$ yields the asymptotics $C\sim\frac{2M_{\mathrm{ADM}}}{r}$ and $\alpha=\sqrt{1-\frac{2M_{\mathrm{ADM}}}{r}}=1-\frac{M_{\mathrm{ADM}}}{r}+\mathcal{O}(r^{-2})$, the latter being the lapse for the exterior Schwarszchild metric and where $M_{\mathrm{ADM}}$ represents the time-conserved Arnowitt-Deser-Misner mass. Finally, demanding $\frac{\partial}{\partial r}\left(r\phi\right)=0$ yields $\phi\sim\frac{\mathrm{const.}}{r}$, the expected dispersal of a massless scalar. These conditions guarantee that the solution given by the network approaches the Schwarzschild asymptotics of an asymptotically flat spacetime, tying the falloff of $\alpha$ and $C$ directly to the ADM mass.
Regarding the resolution of PDEs, the final term in Eq. (22) consists of several addends, each representing the squared modulus of a given equation. On the other hand, the set $\{\lambda_{i}\}$ represents the mixture used in the set of equations. In addition, $\lambda_{\mathcal{IC}}$ denotes the coefficient associated with the initial conditions over time, while the pair $(\lambda_{\mathcal{BC},\>\mathrm{left}},\lambda_{\mathcal{BC},\>\mathrm{right}})$ specifies the weights assigned to the boundary conditions at $r=0$ and $r=R$, respectively.
@@
$\mathcal{L}=\frac{1}{N_{t}N_{r}}\sum_{n=0}^{N_{t}}\omega_{t}\left(t_{n}\right)\sum_{i=0}^{N_{r}}\omega_{r}\left(t_{n},r_{i}\right)\>\mathcal{L}\left(t_{n},r_{i}\right).$ (24)
$\omega_{t}(t_{n})=\mathrm{exp}\left(-\epsilon_{t}\sum_{m=0}^{n-1}\mathcal{L}\left(t_{m}\right)\right).$ (25)
This weight is computed using an exponential function of the cumulative sum of all loss values evaluated at earlier time steps, $\mathcal{L}\left(t_{m<n}\right)$, modulated by a hyperparameter $\epsilon_{t}$, which governs the strength of this causality enforcement mechanism. Within the framework of the original method, the time-dependent loss function $\mathcal{L}(t_{n})$ is obtained by averaging (or, alternatively, summing) over the spatial dimension. In our approach, we propose an analogous strategy for handling the radial dimension by introducing an additional weighting factor $\omega_{r}(t_{n},r_{i})\in[0,1]$,
$\omega_{r}(t_{n},r_{i})=\mathrm{exp}\left(-\epsilon_{r}\sum_{j=0}^{i-1}\mathcal{L}\left(t_{n},r_{j}\right)\right).$ (26)
This weight allows sequential learning along the spatial coordinate for a fixed temporal slice, progressively unlocking spatial points in an iterative manner as the previous losses $\mathcal{L}\left(t_{n},r_{j<i}\right)$ are minimized. Similarly, a new hyperparameter $\epsilon_{r}$ is introduced to regulate the strength of this spatial enforcement mechanism. In general, $\epsilon_{r}$ will be different from its temporal counterpart $\epsilon_{t}$, allowing independent treatment of the learning dynamics in both dimensions.
The use of a loss function of the form (24)-(26) enables the network to carry on a sequential learning, mimicking the behavior of traditional FDM. In this way, learning unfolds progressively through the training process, gradually opening the learning space to the model and preventing the network from “getting ahead” by trying to learn regions that are posterior to those that have not yet been properly minimized. However, despite its clear advantages during training, enforcing sequentiality in learning can also present some drawbacks, including increased computational time and potential issues such as the cost stuck in a specific subset of points $(t,r)$ if the hyperparameter pair $(\epsilon_{t},\epsilon_{r})$ is set too high. Therefore, a thorough sensitivity analysis of the network is necessary with respect to these parameters.
@@ 3.2 Adaptive spacetime sampling
where the loss is calculated using Eq. (22).
@@ 3.3 Model architecture
Figure 2: The proposed methodology for the ModPINN. The embedding block of the model modulates $(t,r)$ via a hyperbolic-tangent “warp”, augments them with low-order polynomials of the normalized coordinates together with a dictionary of $M$ Gaussian transforms, and projects the concatenated features through stacked quadratic (QRes) layers to output the physical set $(\phi,\alpha,C)$.
Throughout this work, several novel methodologies recently introduced in the literature have been considered with the aim of establishing a comparative analysis among them. At the same time, an attempt has been made to introduce a slightly adapted approach of our own, tailored to the specific physical problem under study. This neural architecture is illustrated in the diagram of figure 2, and, for the sake of simplicity in terminology, we shall refer to it as ModPINN—as an alias for “modified PINN”—through the remainder of the text. In general terms, this and any other physics-informed model presented in this work take as input the spacetime coordinates $(t,r)$, process them according to the chosen architecture, and output the set of physical variables, $\left(\phi,\alpha,C\right)_{\Theta}$. These variables are then used to compute the PDEs and, consequently, the loss metric. Although most of the physics of the problem is encapsulated in the said proposed cost function (24), the neural architecture has also been shown to be a fundamental component in terms of performance, as it can usually also be informed by physical principles. Consequently, it is worth highlighting the main aspects of our proposed methodology.
@@
    i.e. an affine transformation, $ax+b$, augmented with a superposition of $K$ independent bump-like deformations (scaled and shifted hyperbolic tangents). Because all parameters $(a,b,c_{k},s_{k},\nu_{k})$ are trainable, each bump introduces a smooth local distortion around its center $\nu_{k}$. By combining multiple bumps, the transformation can selectively stretch or compress different regions of the coordinate axes. Here, $d\in\{t,r\}$ refers to the coordinate dimension and $\mathbf{x}$ to the coordinate tensor. This transformation is particularly valuable for solving critical scenarios in which a black hole forms from gravitational collapse and both sharp and smooth features coexist within the same domain. By adapting the effective resolution of the coordinates, it yields a smooth, infinitely differentiable, and fully adaptive representation in regions with sharp gradients, especially close to the origin. Beyond these transformations, the normalized coordinates are also used to generate some low-order polynomials such as $\tilde{t}^{2}$, $\tilde{r}^{2}$ or $\tilde{t}\>\tilde{r}$, which are concatenated to the previous features, thereby giving more information to the model, which may help overcome spectral bias [Shaghayegh_2025].
  • Lower branch. Simultaneously, before normalizing the coordinates, a dictionary of Gaussian radial basis functions $\{\varphi_{i}(t,r)\}_{i=1}^{M}$ is constructed. This dictionary comprises $M$ different individual functions, each defined as
@@
    and centered at a learnable location $\mu_{i}=\left(\mu_{i,t},\mu_{i,r}\right)$ with also learnable anisotropic scales $l_{i}>0$ for both time and space. Because all parameters are trainable, the model learns where and how wide to place these Gaussian characteristics according to the physical cost function. By stacking $M$ of such functions, the embedding yields a nonlinear, high-dimensional, and smooth representation of the original input space. This enriched map captures localized structure and facilitates the detection of regions with sharp gradients, such as those taking place near formation of black holes.
@@
previous to nonlinear activation, introducing two sets of trainable weights, $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$. Therefore, for problems where linear layers suffice to capture the functional complexity of the target function, one would expect $\mathbf{W}_{2}\to 0$. Conversely, when the model takes advantage from the second-order components, these weights will be non-negligible. Here, $\circ$” stands for the element-wise matrix multiplication (a.k.a. Hadamard product).
@@
Beyond constraining physics to the admissible domain of the problem at hand, activation functions also play a fundamental methodological role. Their choice can be critical to the performance of DL models and, in particular, of PINNs when solving physical scenarios. The influence of activation functions on model performance has attracted considerable attention in the recent literature. Some studies [Maczuga_2023] analyze their effects on relatively simple problems, while others [Abbasi_2024] design problem-specific activations tailored to the physics; an approach that, while potentially effective, may introduce excessive dependence on the problem. A widely shared conclusion is that no single “default” activation function fits all the problems in PINNs, outperforming all others in different fields. Consequently, the choice of activation is highly problem-specific, as some authors [Honghui_Wang_2023] propose to employ a set of basis activation functions, allowing the model itself to select the certain subset from which it can benefit the most.
In our setting, because the weight mappings of the hidden layers are quadratic, it is crucial to employ activation functions that act in a bounded range. Using unbounded activations such as SiLU or ReLU could introduce numerical instabilities by producing arbitrarily large values, particularly during the early stages of training. As noted before, a main concern in the field of PINN methodologies is mitigating their spectral bias, which often causes these models to be unable to resolve problems featuring rapid oscillations and sharp gradients, which is indeed something that we will encounter when solving gravitational collapses. Among the available strategies to counteract this effect, one is to provide the model an additional degree of freedom by including trainable activation functions that the network can adapt to its needs (see, e.g., [Afrah_2025] and references therein). These functions can be made trainable by introducing values that the network learns, with the specific parameterization depending on both the chosen function and the problem at hand. The parameters may be global, layer-specific, or even neuron-specific. In our case, the trainable activation for the hidden layers will be a hyperbolic tangent,
@@ 4 Results
Problem setup. As noted in the introduction, the seminal findings and conclusions of [Choptuik_seminal] are remarkable for their generality and universality. The study examines several families of initial conditions for the scalar field and concludes, for instance, that the exponent $\gamma$ in the black hole mass law (3) appears to be a universal parameter, independent of the initial field configuration. Instead, its dependence is directly related to the deviation of a control parameter $p$ from its critical value $p^{*}$. Despite the well-established universality of criticality that is derived in the gravitational collapse, in our experiments we fix the initial field profile to the “family (a)” (a.k.a. $\phi_{0}(r)$ from now on) configuration from [Choptuik_seminal], expressed as
@@
This profile constitutes the only prior information that is available before any numerical solution—whether via DL or FDMs—is done. However, the initial profiles for the other variables, $\alpha(t,r)$ and $C(t,r)$, are also required for the PINN to learn. These are calculated by setting $t=0$ in the field equations (57) and solving numerically as a preprocessing step. It should be noted that, once the initial pulse is specified, the spatial derivative of the field, $\frac{\partial\phi}{\partial r}$, is straightforwardly determined, while its temporal derivative is not. Because the field evolution equations (5) are second order, the latter must be imposed. Several choices are possible; for example, $\frac{\partial\phi}{\partial t}=0$ (time-symmetric initial data, $\Pi=0$), or $\Pi=\Phi$ (purely ingoing field). In this work, we adopt $\Pi=0$ for all problems and numerical solutions considered. This choice is purely pragmatic: for mathematical simplicity—both in the DL model and in generating the numerical solution—we set $\Pi=0$ without entailing any loss of generality for the study. Regarding the boundary conditions, we impose $C(t,0)=0$ (the so-called elementary flatness at the origin). For $\alpha(t,0)$, there is gauge freedom: $\alpha(t,0)$ may be chosen arbitrarily and $\alpha(t,r)$ can be rescaled by a constant, so that at $r=R$ one may enforce $\alpha(t,R)=\frac{1}{\alpha(t,R)}$. In our setting, we adopt derivative-based boundary conditions (23) instead of Dirichlet conditions, since fixing these variables to constant values might introduce problem-specific dependence. Moreover, since our main objective is to resolve criticality in gravitational collapse using DL-based methods, the chosen scenario is sufficient: it exhibits the relevant interesting behavior and we have a reasonably accurate estimate of the critical parameter, enabling us to distinguish subcritical from supercritical regimes. This setup therefore supports a comprehensive study of the problem, despite the fact that the space of admissible initial field profiles is theoretically unlimited.
@@
  • Loss function. Regarding the loss computation, we set the weights in Eq. (22) as follows: $\lambda_{\mathcal{IC}}=10^{3}$ for the initial conditions; $\lambda_{\mathcal{BC},\mathrm{left}}=\lambda_{\mathcal{BC},\mathrm{right}}=1$ for the boundary conditions at $r=0$ and $r=R$; and $\lambda_{i}=1$ for all the field equations, that is, the PDE residuals. Although several studies conclude that dynamically adapting these weights to the loss of each term can be beneficial, we opt to avoid unnecessary additional complexity and keep them fixed, with $\lambda_{\mathcal{IC}}\gg\{\lambda_{i}\}$ to ensure accurate learning of the initial conditions, which largely determine the physical solution. Regarding the causality and spatial enforcement terms, we use $(\epsilon_{t},\epsilon_{r})=(50,10)$. These values are neither excessively large nor particularly small, leading to a necessary, though nontrivial modification of the vanilla loss function.
  • Spacetime sampling. With respect to the sampled spacetime domain $(t,r)$, several configurations may be considered. In general, the number of points does not need to be large; a standard grid of $N_{t}\times N_{r}=128\times 128$ is typically sufficient. Lighter $(64\times 64)$ and heavier $(256\times 256)$ grids are also examined, particularly to assess how the sampling density affects the model performance. For adaptivity and point movement, we adopt a modest value $\Lambda=1$ in (28), which allows light repositioning of the points.
@@
Numerical comparisons. The field equations (5-7) are straightforwardly solvable in subcritical regimes, i.e., in the absence of black holes, where a vanilla PINN is typically enough. However, as the control parameter $p$ (here, the initial amplitude $\phi_{0}$) approaches its critical value $p^{*}$, the solution develops increasingly sharp and localized gradients (in both space and time). This regime demands a more exhaustive analysis and a modeling specifically tailored to handle such extreme behaviors. Given the inherent difficulty of modeling gravitational collapses, we refrain from relying on a single model. Instead, we compare solutions produced by several PINN-based approaches proposed in recent literature, encompassing methodologies such as the KANs [kan_2025], the PirateNets [piratenets_2024] or the PINN “FP64” [chenhui_2025] which refers to a vanilla PINN where double-precision floating-point format has been considered. Other methodologies—such as PINNsFormer [pinnsformer_2024], which is based on attention, and the recent PINNMamba [mamba_2025], which introduces sub-sequence modeling via State Space Models (SSMs)—were also considered but excluded from the final study due to their training and computational requirements being substantially higher while yielding performances similar to, or even worse than, other simpler models. All training was performed in PyTorch on an NVIDIA A100 (40GB) GPU.
The final results for all models considered and four distinct initial-condition cases are reported in Table 1. From previous numerical studies [Choptuik_seminal, Choptuik_1998], we know with high accuracy that the criticality of the parameter $\phi_{0}$ in the proposed problem occurs near $\phi_{0}=0.1125$, which we adopt as the critical scenario of gravitational collapse. Consequently, we examine two subcritical settings ($\phi_{0}=0.1,0.10625$) and one supercritical setting ($\phi_{0}=0.125$). The exact same loss function, with parameters $(\epsilon_{t},\epsilon_{r})=(50,10)$, is used for all models, as is the adaptive spacetime sapling prescribed by our methodology with $\Lambda=1$. Therefore, the purpose has been to compare neural architectures to handle the gravitational collapse while keeping all other factors fixed. Regarding model-specific settings, we note that, where feasible, all architectures employ six hidden layers with 50 neurons each—except PirateNet, which is organized in blocks rather than layers, and KAN, which uses three layers with five nodes each. Where possible, the hyperbolic tangent is used as an activation function for the hidden layers. Finally, for the proposed ModPINN, the previously introduced standard configuration is adopted.
@@
For each case and each model, we have conducted 10 independent training runs, randomizing the seed—and thus the initial parameters—in every instance, thereby generating some statistics. We report the mean performance alongside the standard deviation. Inspecting the results, we observe that the critical and supercritical cases exhibit the largest variability across most models, highlighting the difficulty in solving these scenarios. For the critical case, ModPINN achieves the best performance in both mean accuracy and variability, indicating greater stability than the alternatives for this scenario. The same conclusion holds for the supercritical case with $\phi_{0}=0.125$. In contrast, in the two subcritical settings, the purely quadratic PINN (QRes) leads to the strongest results; however, as noted, these cases are easily solved by a vanilla PINN due to their relative simplicity and smoothness, so the outcome and the similarity across models are generally accepted. It is therefore excepted that QRes ranks among the strongest models: quadratic-weight layers help mitigate the frequency bias typical of PINNs. These layers also aid in resolving sharp localized gradients, as evidenced by the proposed ModPINN. However, for $\phi_{0}=0.1125$ and $0.125$, QRes is not among the top performers. The principal reason is that, while quadratic layers can substantially benefit such problems, they may also induce numerical instabilities. Therefore, implementing the proposed embedding—which normalizes input coordinates to numerically manageable values and generates additional features—indeed improves performance on criticality and gravitational collapse cases. These are precisely the regimes of greatest interest as they reveal the major physical phenomena.

### v2

@@ 7. Define uniform target masses: For each $i$,
In our methodology, a two-dimensional spacetime domain is numerically discretized as $(t,r)\in[0,T]\times[0,R]$, where the parameters $T$ and $R$ denote the maximum values reached in the numerical simulation and must be chosen sufficiently large to encompass all relevant physical information. The resulting grid thus comprises $N_{\mathrm{t}}$ points in the temporal dimension and $N_{\mathrm{r}}$ in the radial direction. Therefore, the original loss function depends on both time and space, since a loss value is computed at each collocation point. In that sense, the first three terms in the loss equation
@@
The enforcement of restrictions $\left|\mathcal{R}_{\alpha}\right|^{2},\left|\mathcal{R}_{C}\right|^{2}\to 0$ yields the asymptotics $C\sim\frac{2M_{\mathrm{ADM}}}{r}$ and $\alpha=\sqrt{1-\frac{2M_{\mathrm{ADM}}}{r}}=1-\frac{M_{\mathrm{ADM}}}{r}+\mathcal{O}(r^{-2})$, the latter being the lapse for the exterior Schwarzschild metric and where $M_{\mathrm{ADM}}$ represents the time-conserved Arnowitt-Deser-Misner mass. Finally, demanding $\frac{\partial}{\partial r}\left(r\phi\right)=0$ yields $\phi\sim\frac{\mathrm{const.}}{r}$, the expected dispersal of a massless scalar. These conditions guarantee that the solution given by the network approaches the Schwarzschild asymptotics of an asymptotically flat spacetime, tying the falloff of $\alpha$ and $C$ directly to the ADM mass.
Regarding the resolution of PDEs, the final term in (22) consists of several terms, each representing the squared modulus of a given equation. On the other hand, the set $\{\lambda_{i}\}$ represents the weights assigned to the different equations. In addition, $\lambda_{\mathcal{IC}}$ denotes the coefficient associated with the initial conditions over time, while the pair $(\lambda_{\mathcal{BC},\>\mathrm{left}},\lambda_{\mathcal{BC},\>\mathrm{right}})$ specifies the weights assigned to the boundary conditions at $r=0$ and $r=R$, respectively.
@@
$\mathcal{L}=\frac{1}{N_{\mathrm{t}}N_{\mathrm{r}}}\sum_{n=0}^{N_{\mathrm{t}}}\omega_{t}\left(t_{n}\right)\sum_{i=0}^{N_{\mathrm{r}}}\omega_{r}\left(t_{n},r_{i}\right)\>\mathcal{L}\left(t_{n},r_{i}\right).$ (24)
$\omega_{t}(t_{n})=\mathrm{exp}\left(-\epsilon_{\mathrm{t}}\sum_{m=0}^{n-1}\mathcal{L}\left(t_{m}\right)\right).$ (25)
This weight is computed using an exponential function of the cumulative sum of all loss values evaluated at earlier time steps, $\mathcal{L}\left(t_{m<n}\right)$, modulated by a hyperparameter $\epsilon_{\mathrm{t}}$, which governs the strength of this causality enforcement mechanism. Within the framework of the original method, the time-dependent loss function $\mathcal{L}(t_{n})$ is obtained by averaging (or, alternatively, summing) over the spatial dimension. In our approach, we propose an analogous strategy for handling the radial dimension by introducing an additional weighting factor $\omega_{r}(t_{n},r_{i})\in[0,1]$,
$\omega_{r}(t_{n},r_{i})=\mathrm{exp}\left(-\epsilon_{\mathrm{r}}\sum_{j=0}^{i-1}\mathcal{L}\left(t_{n},r_{j}\right)\right).$ (26)
This weight allows sequential learning along the spatial coordinate for a fixed temporal slice, progressively unlocking spatial points in an iterative manner as the previous losses $\mathcal{L}\left(t_{n},r_{j<i}\right)$ are minimized. Similarly, a new hyperparameter $\epsilon_{\mathrm{r}}$ is introduced to regulate the strength of this spatial enforcement mechanism. In general, $\epsilon_{\mathrm{r}}$ will be different from its temporal counterpart $\epsilon_{\mathrm{t}}$, allowing independent treatment of the learning dynamics in both dimensions.
The use of a loss function of the form (24)-(26) enables the network to perform sequential learning, mimicking the behavior of traditional FDM. In this way, learning unfolds progressively through the training process, gradually opening the learning space to the model and preventing the network from “getting ahead” by trying to learn regions that are posterior to those that have not yet been properly minimized. However, despite its clear advantages during training, enforcing sequentiality in learning can also present some drawbacks, including increased computational time and potential issues such as the cost gets stuck in a specific subset of points $(t,r)$ if the hyperparameter pair $(\epsilon_{\mathrm{t}},\epsilon_{\mathrm{r}})$ is set too high. Therefore, a thorough sensitivity analysis of the network is necessary with respect to these parameters.
@@ 3.2 Adaptive spacetime sampling
where the loss is calculated using (22).
@@ 3.3 Model architecture
Figure 2: The proposed methodology for the ModPINN. The embedding block of the model modulates $(t,r)$ via a hyperbolic-tangent “warp”, augments them with low-order polynomials of the normalized coordinates together with a dictionary of $\mathrm{M}$ Gaussian transforms, and projects the concatenated features through stacked quadratic (QRes) layers to output the physical set $(\phi,\alpha,C)$.
Throughout this work, several novel methodologies recently introduced in the literature have been considered with the aim of establishing a comparative analysis among them. At the same time, an attempt has been made to introduce a slightly adapted approach of our own, tailored to the specific physical problem under study. This neural architecture is illustrated in the diagram of figure 2, and, for the sake of simplicity in terminology, we shall refer to it as ModPINN—as an alias for “modified PINN”—throughout the remainder of the text. In general terms, this and any other physics-informed model presented in this work take as input the spacetime coordinates $(t,r)$, process them according to the chosen architecture, and output the set of physical variables, $\left(\phi,\alpha,C\right)_{\Theta}$. These variables are then used to compute the PDEs and, consequently, the loss metric. Although most of the physics of the problem is encapsulated in the proposed cost function (24), the neural architecture has also been shown to be a fundamental component in terms of performance, as it can usually be informed by physical principles as well. Consequently, it is worth highlighting the main aspects of our proposed methodology.
@@
    i.e. an affine transformation, $ax+b$, augmented with a superposition of $\mathrm{K}$ independent bump-like deformations (scaled and shifted hyperbolic tangents). Because all parameters $(a,b,c_{k},s_{k},\nu_{k})$ are trainable, each bump introduces a smooth local distortion around its center $\nu_{k}$. By combining multiple bumps, the transformation can selectively stretch or compress different regions of the coordinate axes. Here, $d\in\{t,r\}$ refers to the coordinate dimension and $\mathbf{x}$ to the coordinate tensor. This transformation is particularly valuable for solving critical scenarios in which a black hole forms from gravitational collapse and both sharp and smooth features coexist within the same domain. By adapting the effective resolution of the coordinates, this block yields a smooth, infinitely differentiable, and fully adaptive representation in regions with sharp gradients, especially close to the origin. Beyond these transformations, the normalized coordinates are also used to generate some low-order polynomials such as $\tilde{t}^{2}$, $\tilde{r}^{2}$ or $\tilde{t}\>\tilde{r}$, which are concatenated to the previous features, thereby giving more information to the model, which may help overcome spectral bias [Shaghayegh_2025].
  • Lower branch. Simultaneously, before normalizing the coordinates, a dictionary of Gaussian radial basis functions $\{\varphi_{i}(t,r)\}_{i=1}^{M}$ is constructed. This dictionary comprises $\mathrm{M}$ different individual functions, each defined as
@@
    and centered at a learnable location $\mu_{i}=\left(\mu_{i,t},\mu_{i,r}\right)$ with learnable anisotropic scales $l_{i}>0$ for both time and space. Because all parameters are trainable, the model learns where and how wide to place these Gaussian characteristics according to the physical cost function. By stacking $\mathrm{M}$ of such functions, the embedding yields a nonlinear, high-dimensional, and smooth representation of the original input space. This enriched map captures localized structure and facilitates the detection of regions with sharp gradients, such as those taking place near formation of black holes.
@@
prior to nonlinear activation, introducing two sets of trainable weights, $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$. Therefore, for problems where linear layers suffice to capture the functional complexity of the target function, one would expect $\mathbf{W}_{2}\to 0$. Conversely, when the model takes advantage of the second-order components, these weights will be non-negligible. Here, $\circ$” stands for the element-wise matrix multiplication (a.k.a. Hadamard product).
@@
Beyond constraining physics to the admissible domain of the problem at hand, activation functions also play a fundamental methodological role. Their choice can be critical to the performance of DL models and, in particular, of PINNs when solving physical scenarios. The influence of activation functions on model performance has attracted considerable attention in the recent literature. Some studies [Maczuga_2023] analyze their effects on relatively simple problems, while others [Abbasi_2024] design problem-specific activations tailored to the physics; an approach that, while potentially effective, may introduce excessive dependence on the problem. A widely shared conclusion is that no single “default” activation function fits all the problems in PINNs, outperforming all others in different fields. Consequently, the choice of activation is highly problem-specific, as some authors [Honghui_Wang_2023] propose employing a set of basis activation functions, allowing the model itself to select the certain subset from which it can benefit the most.
In our setting, because the weight mappings of the hidden layers are quadratic, it is crucial to employ activation functions that act in a bounded range. Using unbounded activations such as SiLU or ReLU could introduce numerical instabilities by producing arbitrarily large values, particularly during the early stages of training. As noted before, a main concern in the field of PINN methodologies is mitigating their spectral bias, which often causes these models to be unable to resolve problems featuring rapid oscillations and sharp gradients, which is indeed something that we will encounter when solving gravitational collapses. Among the available strategies to counteract this effect, one is to provide the model with an additional degree of freedom by including trainable activation functions that the network can adapt to its needs (see, e.g., [Afrah_2025] and references therein). These functions can be made trainable by introducing values that the network learns, with the specific parameterization depending on both the chosen function and the problem at hand. The parameters may be global, layer-specific, or even neuron-specific. In our case, the trainable activation for the hidden layers will be a hyperbolic tangent,
@@ 4 Results
Problem setup. As noted in the introduction, the seminal findings and conclusions of [Choptuik_seminal] are remarkable for their generality and universality. The study examines several families of initial conditions for the scalar field and concludes, for instance, that the exponent $\gamma$ in the black hole mass law (3) appears to be a universal quantity, independent of the initial field configuration. Instead, its dependence is directly related to the deviation of a control parameter $p$ from its critical value $p^{*}$. Despite the well-established universality of critical phenomena in gravitational collapse, in our experiments we fix the initial field profile to the “family (a)” (a.k.a. $\phi_{0}(r)$ from now on) configuration from [Choptuik_seminal], expressed as
@@
This profile constitutes the only prior information that is available before any numerical solution—whether via DL or FDMs—is obtained. However, the initial profiles for the other variables, $\alpha(t,r)$ and $C(t,r)$, are also required for the PINN to learn. These are calculated by setting $t=0$ in the field equations (57) and solving numerically as a preprocessing step. It should be noted that, once the initial pulse is specified, the spatial derivative of the field, $\frac{\partial\phi}{\partial r}$, is straightforwardly determined, while its temporal derivative is not. Because the field evolution equations (5) are second order, the latter must be imposed. Several choices are possible; for example, $\frac{\partial\phi}{\partial t}=0$ (time-symmetric initial data, $\Pi=0$), or $\Pi=\Phi$ (purely ingoing field). In this work, we adopt $\Pi=0$ for all problems and numerical solutions considered. This choice is purely pragmatic: for mathematical simplicity—both in the DL model and in generating the numerical solution—we set $\Pi=0$ without entailing any loss of generality for the study. Regarding the boundary conditions, we impose $C(t,0)=0$ (the so-called elementary flatness at the origin). For $\alpha(t,0)$, there is gauge freedom: $\alpha(t,0)$ may be chosen arbitrarily and $\alpha(t,r)$ can be rescaled by a constant, so that at $r=R$ one may enforce $\alpha(t,R)=\frac{1}{\alpha(t,R)}$. In our setting, we adopt derivative-based boundary conditions (23) instead of Dirichlet conditions, since fixing these variables to constant values might introduce problem-specific dependence. Moreover, since our main objective is to resolve criticality in gravitational collapse using DL-based methods, the chosen scenario is sufficient: it exhibits the relevant interesting behavior and we have a reasonably accurate estimate of the critical parameter, enabling us to distinguish subcritical from supercritical regimes. This setup therefore supports a comprehensive study of the problem, despite the fact that the space of admissible initial field profiles is theoretically unlimited.
@@
  • Loss function. Regarding the loss computation, we set the weights in (22) as follows: $\lambda_{\mathcal{IC}}=10^{3}$ for the initial conditions; $\lambda_{\mathcal{BC},\mathrm{left}}=\lambda_{\mathcal{BC},\mathrm{right}}=1$ for the boundary conditions at $r=0$ and $r=R$; and $\lambda_{i}=1$ for all the field equations, that is, the PDE residuals. Although several studies conclude that dynamically adapting these weights to the loss of each term can be beneficial, we opt to avoid unnecessary additional complexity and keep them fixed, with $\lambda_{\mathcal{IC}}\gg\{\lambda_{i}\}$ to ensure accurate learning of the initial conditions, which largely determine the physical solution. Regarding the causality and spatial enforcement terms, we use $(\epsilon_{\mathrm{t}},\epsilon_{\mathrm{r}})=(50,10)$. These values are neither excessively large nor particularly small, leading to a necessary, though nontrivial modification of the vanilla loss function.
  • Spacetime sampling. With respect to the sampled spacetime domain $(t,r)$, several configurations may be considered. In general, the number of points does not need to be large; a standard grid of $N_{\mathrm{t}}\times N_{\mathrm{r}}=128\times 128$ is typically sufficient. Lighter $(64\times 64)$ and heavier $(256\times 256)$ grids are also examined, particularly to assess how the sampling density affects the model performance. For adaptivity and point movement, we adopt a modest value $\Lambda=1$ in (28), which allows light repositioning of the points.
@@
Numerical comparisons. The field equations (5-7) are straightforwardly solvable in subcritical regimes, i.e., in the absence of black holes, where a vanilla PINN is typically enough. However, as the control parameter $p$ (here, the initial amplitude $\phi_{0}$) approaches its critical value $p^{*}$, the solution develops increasingly sharp and localized gradients (in both space and time). This regime demands a more exhaustive analysis and a modeling specifically tailored to handle such extreme behaviors. Given the inherent difficulty of modeling gravitational collapses, we refrain from relying on a single model. Instead, we compare solutions produced by several PINN-based approaches proposed in recent literature, encompassing methodologies such as the KANs [kan_2025], the PirateNets [piratenets_2024] or the PINN “FP64” [chenhui_2025] which refers to a vanilla PINN where double-precision floating-point format has been considered. Other methodologies—such as PINNsFormer [pinnsformer_2024], which is based on attention, and the recent PINNMamba [mamba_2025], which introduces subsequence modeling via State Space Models (SSMs)—were also considered but excluded from the final study due to their training and computational requirements being substantially higher while yielding performances similar to, or even worse than, other simpler models. All training was performed in PyTorch on an NVIDIA A100 (40GB) GPU.
The final results for all models considered and four distinct initial-condition cases are reported in Table 1. From previous numerical studies [Choptuik_seminal, Choptuik_1998], we know with high accuracy that the critical value of the parameter $\phi_{0}$ in the proposed problem occurs near $\phi_{0}=0.1125$, which we adopt as the critical scenario of gravitational collapse. Consequently, we examine two subcritical settings ($\phi_{0}=0.1,0.10625$) and one supercritical setting ($\phi_{0}=0.125$). The exact same loss function, with parameters $(\epsilon_{\mathrm{t}},\epsilon_{\mathrm{r}})=(50,10)$, is used for all models, as is the adaptive spacetime sampling prescribed by our methodology with $\Lambda=1$. Therefore, the purpose has been to compare neural architectures to handle the gravitational collapse while keeping all other factors fixed. Regarding model-specific settings, we note that, where feasible, all architectures employ six hidden layers with 50 neurons each—except PirateNet, which is organized in blocks rather than layers, and KAN, which uses three layers with five nodes each. Where possible, the hyperbolic tangent is used as an activation function for the hidden layers. Finally, for the proposed ModPINN, the previously introduced standard configuration is adopted.
@@
For each case and each model, we have conducted 10 independent training runs, randomizing the seed—and thus the initial parameters—in every instance, thereby generating some statistics. We report the mean performance alongside the standard deviation. Inspecting the results, we observe that the critical and supercritical cases exhibit the largest variability across most models, highlighting the difficulty in solving these scenarios. For the critical case, ModPINN achieves the best performance in both mean accuracy and variability, indicating greater stability than the alternatives for this scenario. The same conclusion holds for the supercritical case with $\phi_{0}=0.125$. In contrast, in the two subcritical settings, the purely quadratic PINN (QRes) leads to the strongest results; however, as noted, these cases are easily solved by a vanilla PINN due to their relative simplicity and smoothness, so the outcome and the similarity across models are generally accepted. It is therefore expected that QRes ranks among the strongest models: quadratic-weight layers help mitigate the frequency bias typical of PINNs. These layers also aid in resolving sharp localized gradients, as evidenced by the proposed ModPINN. However, for $\phi_{0}=0.1125$ and $0.125$, QRes is not among the top performers. The principal reason is that, while quadratic layers can substantially benefit such problems, they may also induce numerical instabilities. Therefore, implementing the proposed embedding—which normalizes input coordinates to numerically manageable values and generates additional features—indeed improves performance on criticality and gravitational collapse cases. These are precisely the regimes of greatest interest as they reveal the major physical phenomena.